OpenStudy (anonymous):

Charlotte has two job offers as a car sales person. The first job will pay her $600 a month plus 2% on the price of her car sales. The second job will pay $1000 a month plus 1% on the price of her sales. How much must she sell each month to make the same income from the two jobs? A. $20,000 B. $30,000 C. $40,000 D. $50,000

OpenStudy (anonymous):

First job: pay = 0.02x + 600 Second job: pay = 0.01x +1000 Where x represents the amount of her sales... to see when her pay is the same set the equations equal to each other. 0.02x + 600 = 0.01x + 1000 Solve. 0.01x = 400 x= 40 000 She would need have $40,000 in sales for her pay to be the same.